# Leveraging Hierarchical Taxonomies in Prompt-based Continual Learning

### v2

@@ 5.2 Experimental result
Ours (TCL) 94.52 ±$\pm 0.22$ 1.02 ±$\pm 0.18$ 77.01 ±$\pm 0.12$ 4.03 ±$\pm 0.25$ 95.35 ±$\pm 0.18$ 0.20 ±$\pm 0.16$ 88.33 ±$\pm 0.18$ 1.98 ±$\pm 0.22$
Our training strategy improves model performance significantly.
Figure 7 reports the ablation studies demonstrating the effectiveness of our training strategy. Particularly, compared to training tasks independently using Cross Entropy loss $\mathcal{L}_{CE}$ like in DualP, L2P, and CODA-P, exploiting the relationships between data classes with the label-based hierarchical taxonomy and the WD-based cost matrix helps improve FAA by about 5% to 10% (Figure 7(a)). Besides, when examining the role of exploiting additional prior information from pretrained backbones using the OT approach, we see that FAA is improved from 0.6% to 0.8% (Figure 7(b)). These results demonstrate the positive impact of this component, confirming the importance of exploiting correlations between class data during training. In both figures, the improvements on Split-CIFAR100 and 5-Datasets are the lowest, while it is more pronounced on Split-CUB-200. This may be because the groups of these two datasets (Split-CIFAR100 and 5-Datasets) have fewer overlapping classes, as the classes in each group likely have more recognizable features. Meanwhile, Split-CUB-200 is a dataset about birds, with images that can be difficult for human eyes to recognize, thus so our method performs better.
In addition, Figure 7(c) provides the experimental results on Split-CUB-200 dataset, when varying $\alpha$ and $\beta$, which control the intensification of impact on each leaf group of $\mathcal{L_{G}}$ during training. The data shows that with a large enough value of $\beta$, our HiT is not sensitive to $\alpha$ within its acceptable range. Conversely, if $\alpha$ is small, the quality of the model can change more significantly.
Furthermore, Figure 4 illustrates the effect of our method in improving model’s representation learning on Split-CIFAR100. Specifically, the classes are better clustered, and the separation between them is more distinct. Especially, the classes ’oak tree,’ ’willow tree,’ and ’pine tree’ are divided into clear clusters, rather than being mixed together as in the traditional training strategy, where tasks are trained independently.
@@
[…]
Figure 7: Ablation study about our training strategy.
Our testing strategy has positive effects for final prediction
Table 3 illustrates the improvement of FAA on all considered datasets when applying our testing strategy, from about 0.4% to 0.6%. This proves the approach to be effective, as the information from each cluster provides a reference channel that helps determine the identity of the classes, offering a good suggestion for future studies.
Table 3: Effectiveness of our testing strategy.
Dataset Split CIFAR-100 Split Imagenet-R 5-Datasets Split CUB-200
Normal testing 93.94 76.41 95.02 87.93
Our testing strategy 94.52 77.01 95.35 88.33
6 Conclusion and Limitation discussion
@@
In this work, we demonstrate the importance of organizing and exploiting data meaningfully rather than lumping it together for training. Organizing data into a tree-like taxonomy based on label information gives us a new perspective on the data. Particularly, we can divide them into small groups containing the classes that are likely to confuse models. This approach encourages the model to focus and build deeper knowledge for each group, thereby reducing forgetting and motivating more effective learning in subsequent tasks. Additionally, we introduce a new perspective by leveraging the initial behavior of pretrained models, providing an additional information channel to further improve performance. Besides, our testing strategy has shown positive effects by exploiting group knowledge during inference. Finally, experimental results demonstrate the effectiveness of these components and our superiority over state-of-the-art baselines.
Limitation discussion.
@@
Despite this novel perspective, the quality of the hierarchical taxonomy depends on the quality of expert knowledge. For example, if similar image classes are not assigned to the same leaf group in this label-based taxonomy, the constraint we put on each such group may not perform as expected. Furthermore, although the testing strategy shows positive results, to exploit group knowledge more efficiently, it is necessary to further investigate to understand the characteristics and hidden structure of data.
Appendices
@@ A.2 Baselines
In the main paper, we use CL methods with pre-trained ViT as the backbone. We group them into (a) the group using a common prompt pool for all tasks, and (b) the group dedicating distinct prompt sets for each task:
@@
(2) DualPrompt [34]: The prompt-based method aims to address L2P’s limitations by attaching complementary prompts to the pre-trained backbone, rather than only at input. DualP introduces additional prompt sets for each task to leverage task-specific instructions alongside invariant information from the common pool. However, like L2P, it does not focus on efficiently learning the classification head. Additionally, selecting the wrong prompt ID for task-specific instructions during testing can negatively impact model performance.
@@
(6) HiDe-Prompt [30]: a recent SOTA prompt-based method that decomposes learning CIL into 3 modules: a task inference, a within-task predictor and a task-adaptive predictor. The second module trains prompts for each task with a contrastive regularization that tries to push features of new tasks away from prototypes of old ones. To predict task identity, it trains a classification head on top of the pre-trained ViT. TAP is similar to a fine-tuning step that aims to alleviate classifier bias using the Gaussian distribution of all classes seen so far. However, this method does not declare the relationship between data during training, thereby missing the opportunity to improve model performance.
@@ A.4 Implementation Details
Our implementation basically aligns with the methodologies employed in prior research [35, 34, 25]. Specifically, our framework incorporates the use of a pre-trained Vision Transformer (ViT-B/16) as the backbone architecture. For the optimization process, we utilized the Adam optimizer, configured with hyper-parameters $\beta_{1}$ set to $0.9$ and $\beta_{2}$ set to $0.999$. The training process was conducted using batches of $24$ samples, and a fixed learning rate of $0.03$ was applied across all models except for CODA-Prompt. For CODA-Prompt, we employed a cosine decaying learning rate strategy, starting at $0.001$. Additionally, a grid search technique was implemented to determine the most appropriate number of epochs for effective training. Regarding the pre-processing of input data, images were resized to a standard dimension of $224\times 224$ pixels and normalized within a range of $[0,1]$ to ensure consistency in input data format. The detailed values of the parameters can be found in our source code.
@@ Appendix B Using ChatGPT to build tree-like taxonomy during a sequence of tasks incrementally
When the list [$\cdots$’] is ["leopard", "rabbit", "mouse", "camel", "trout", "aquarium_fish", "snake", "rose", "lawn_mower", "bottle"], we can obtain an output as follows:

### v3

@@ 4.3 Harnessing Prior Knowledge from Pretrained Models via Optimal Transport
Ours (TCL) 93.94 ±$\pm 0.23$ 1.05 ±$\pm 0.15$ 76.41 ±$\pm 0.12$ 4.02 ±$\pm 0.22$ 95.02 ±$\pm 0.20$ 0.21 ±$\pm 0.15$ 87.93 ±$\pm 0.22$ 2.01 ±$\pm 0.23$
Firstly, to extract the relationship between the classes, we compute the L2-Wasserstein distance of feature distributions of each class pair. Specifically, we denote $D_{c_{i}}^{\Phi}$ be the distribution of class $c_{i}$ on the latent space of the pretrained model $f_{\Phi}$ to distinguish it from the class distribution on prompted latent space $f_{\Phi,\bm{P}}$, as described in Section 4.2. The distribution $D_{c_{i}}^{\Phi}$ is obtained in the form of a Gaussian Mixture model at the end of the respective task, only once and then kept fixed. In this way, when $t$ tasks have arrived, we have the corresponding sets of distributions $\{D_{c}^{\Phi}\}_{c\in Y_{1,t}}$ of all $m_{t}$ classes so far. Therefore, we gradually complete the WD-based matrix between pairs of classes:
$M=[W_{2}(D_{c_{i}}^{\Phi},D_{c_{j}}^{\Phi})]_{m_{t}\times m_{t}}.$ (2)
The fact is that the smaller the $W_{2}$ between the representations of two classes, the closer they are, and the more force can be needed to separate them. Therefore, we employ these results to obtain the corresponding weight factors, which control how much constraint needs to be applied to each pair. In particular, we compute the weight matrix:
$\Gamma=\left[\gamma_{ij}\right]_{m_{t}\times m_{t}}=\left[1/{\exp(M_{ij}/\delta)}\right]_{m_{t}\times m_{t}}$ (3)
where $\delta$ is a temperature. We then apply this information to obtain a weighted version of $\mathcal{L}_{g}$, in which the closer the two class distributions are, the larger the weight assigned, and they will be focused to push away. Consequently, our regularization loss becomes:
$\mathcal{L}_{g}(\cdot)=-\log\sum_{{\bm{x}}^{\prime}\in X^{g}_{t}|y_{{\bm{x}}^{\prime}}=y_{\bm{x}}}\frac{u({\bm{z}}_{x}\cdot{\bm{z}}_{x^{\prime}})}{\sum_{\bar{x}\in X^{g}_{\overline{1,t}}}\gamma_{y_{x}y_{\tilde{x}}}u({\bm{z}}_{x}\cdot{\bm{z}}_{\bar{x}})}$ (4)
This strategy is completely economical and aligns well with the CL learning scheme, as the distance between representations of each class pair only needs to be computed once, and the matrix $M$ is continuously expanded when new classes arrive. Practically, when learning a new task, the first epoch is for capturing information about the behavior of the pretrained model on the data of this task. Moreover, this approach is also aligned with the findings in Cognitive Science [34, 2], showing that the accumulated experiences from past learning create momentum for learning new skills more effectively.
Finally, our objective function during training can be formulated as follows:
$\mathcal{L}=\mathcal{L}_{CE}+\alpha\mathcal{L}_{g}+\beta\mathcal{L}_{all}$ (5)
5 Experiments
5.1 Experimental setup
Datasets. We use 4 common CIL benchmarks, including Split CIFAR-100, Split ImageNet-R, 5-Datasets, and Split CUB-200 (see Appendix A.1).
Baselines. We compare ours with the latest state-of-the-art prompt-based methods, including the methods using shared prompts for all tasks: L2P [49], DualPrompt [48], OVOR [18]; and the methods dedicates a distinct prompt set for each task like: S-Prompt++ [47], CODA-Prompt [38], HiDe-Prompt [44], CPP [25] (see Appendix A.2).
Metrics. We use two main metrics, including the Final Average Accuracy (FAA) and the Final Forgetting Measure (FFM), denoting the average accuracy and forgetting of all tasks after learning the sequence of tasks, respectively (see Appendix A.2). Besides, the detailed implementation is described in Appendix A.4.
@@ 5.1 Experimental setup
Figure 6: Ablation studies about our training strategy.
5.2 Experimental result
Our approach achieves superior results compared to baselines.
Table 1 presents the overall performance comparison between our proposed method and all the baselines. The key observation is that our method is the strongest one, by the gap between our method and the runner-up method being around 1.5% in terms of FAA on all considered datasets. Additionally, the results show that our method avoids forgetting better than all baselines, by the gap up to 17% on 5-Datasets.
Ablation studies about our training strategy
Figure 6 reports the ablation studies demonstrating the effectiveness of our training strategy. Particularly, compared to training tasks independently using Cross Entropy loss $\mathcal{L}_{CE}$ like in DualP, L2P, and CODA-P, exploiting the relationships between data classes with the label-based hierarchical taxonomy and the WD-based cost matrix helps improve FAA by about 5% to 10% (Figure 6(a)). Besides, when examining the role of exploiting additional prior information from pretrained backbones using the OT approach, we see that FAA is improved from 0.6% to 0.8% (Figure 6(b)). These results demonstrate the positive impact of this component, confirming the importance of exploiting correlations between class data during training. In both figures, the improvements on Split-CIFAR100 and 5-Datasets are the lowest, while it is more pronounced on Split-CUB-200. This may be because the groups of these two datasets (Split-CIFAR100 and 5-Datasets) have fewer overlapping classes, as the classes in each group likely have more recognizable features. Meanwhile, Split-CUB-200 is a dataset of birds with images that can be difficult for human eyes to recognize, thus so our method performs better.
In addition, Figure 6(c) provides the experimental results on the dataset Split-CUB-200, when varying $\alpha$ and $\beta$, which control the intensification of impact on each leaf group of $\mathcal{L_{G}}$ during training. The data shows that with a large enough value of $\beta$, our TCL is not sensitive to $\alpha$ within its acceptable range. Conversely, if $\alpha$ is small, the quality of the model can change more significantly.
t-SNE visualization of representation on latent space
Furthermore, Figure 4 illustrates the effect of our method in improving the model’s representation learning on Split-CIFAR100. Specifically, the classes are better clustered, and the separation between them is more distinct. Especially, the classes ’oak tree,’ ’willow tree,’ and ’pine tree’ are divided into clear clusters, rather than being mixed together as in the traditional training strategy, where tasks are trained independently.
The superiority of our proposed method on various types of pretrained backbones.
Table 2 illustrates that our method with the training strategy only consistently outperforms the strongest baseline (HiDE-Prompt) by the gap from about 0.5% to 1.5% in all cases.
Backbone Split CIFAR-100 Split Imagenet-R
TCL HiDE-Prompt TCL HiDE-Prompt
Sup-21K 93.94 92.61 76.41 75.06
iBOT-21K 94.01 93.02 72.12 70.83
iBOT-1K 94.27 93.48 72.80 71.33
DINO-1K 94.12 93.51 69.25 68.11
MoCo-1K 92.32 91.57 64.23 63.77
Table 2: Comparison when using different pretrained backbones
6 Conclusion and Discussion
@@
In this work, we demonstrate the importance of organizing and exploiting data meaningfully rather than lumping it together for training. In particular, we propose organizing data into a tree-like taxonomy based on their label information, helping to identify groups of class data that are likely to confuse models. In this way, we encourage the models to focus and build deeper knowledge about each group, reducing forgetting and motivating more effective learning in subsequent tasks. Additionally, we propose leveraging the initial behavior of pretrained models to unveil a hidden structure of training data, providing an additional perspective to improve performance further. Finally, experimental results demonstrate the effectiveness of these components and our superiority over state-of-the-art baselines.
Limitations.
@@
Despite this novel approach, the quality of the hierarchical taxonomy depends on the quality of expert knowledge. For example, if similar image classes are not assigned to the same leaf group in this label-based taxonomy, the constraint we put on each such group may not perform as expected.
Supplement to “Leveraging Hierarchical Taxonomies in Prompt-based Continual Learning“
@@ A.2 Baselines
In the main paper, we use CL methods with pretrained ViT as the backbone. We group them into (a) the group using a common prompt pool for all tasks, and (b) the group dedicating distinct prompt sets for each task:
@@
(2) DualPrompt [48]: The prompt-based method aims to address L2P’s limitations by attaching complementary prompts to the pretrained backbone, rather than only at input. DualP introduces additional prompt sets for each task to leverage task-specific instructions alongside invariant information from the common pool. However, like L2P, it does not focus on efficiently learning the classification head. Additionally, selecting the wrong prompt ID for task-specific instructions during testing can negatively impact model performance.
@@
(6) HiDe-Prompt [44]: a recent SOTA prompt-based method that decomposes learning CIL into 3 modules: a task inference, a within-task predictor and a task-adaptive predictor. The second module trains prompts for each task with a contrastive regularization that tries to push features of new tasks away from prototypes of old ones. To predict task identity, it trains a classification head on top of the pretrained ViT. TAP is similar to a fine-tuning step that aims to alleviate classifier bias using the Gaussian distribution of all classes seen so far. However, this method does not declare the relationship between data during training, thereby missing the opportunity to improve model performance.
@@ A.4 Implementation Details
Our implementation basically aligns with the methodologies employed in prior research [49, 48, 38]. Specifically, our framework incorporates the use of a pretrained Vision Transformer (ViT-B/16) as the backbone architecture. For the optimization process, we utilized the Adam optimizer, configured with hyper-parameters $\beta_{1}$ set to $0.9$ and $\beta_{2}$ set to $0.999$. The training process was conducted using batches of $24$ samples, and a fixed learning rate of $0.03$ was applied across all models except for CODA-Prompt. For CODA-Prompt, we employed a cosine decaying learning rate strategy, starting at $0.001$. Additionally, a grid search technique was implemented to determine the most appropriate number of epochs for effective training. Regarding the pre-processing of input data, images were resized to a standard dimension of $224\times 224$ pixels and normalized within a range of $[0,1]$ to ensure consistency in input data format. The detailed values of the parameters can be found in our source code.
@@ Appendix B Using ChatGPT to build tree-like taxonomy during a sequence of tasks incrementally
Example output, when the list [$\cdots$’] is ["leopard", "rabbit", "mouse", "camel", "trout", "aquarium_fish", "snake", "rose", "lawn_mower", "bottle"]: